OpenStudy (anonymous):

Find the x-, y-, and z-intercepts for the equation. 3x – 5y + 10z = 15

OpenStudy (anonymous):

So basically any equation showing x , y , z intercepts is of the form - x/a + y/b + z/c = 1. so can u tell me which number can 15 divide with so that the ans is 1 ?

OpenStudy (anonymous):

15?

OpenStudy (anonymous):

yeah. so u will divide the whole equation by 15. Which becomes - 3x/15 -5y/15 +10z/15 = 15/15. which is x/5 -y/3 + 2z/3 = 1 now to find the intercepts u will evaluate each fraction equal to 1. so x/5 = 1, which gives x = 5, -y/3 = 1 which gives y = -3, and 2z/3 = 1 which gives z = 3/2.